(10 pts) In an advertisement, a pizza shop claims that its mean delivery time is less than 30 minutes. A random sample of 36 delivery imes has a mean of 28.5 minutes and a standard deviation of 3.5 minutes. Is there enough evidence to support the claim at α-.05 ? (a) Set up the null and alternative hypotheses (b) Find the test statistic. (c) Find the rejection region and state your conclusion. (d) What is the P-value for this test?
Answer #1

Solution :

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 30

Ha : \mu   \neq 30

\bar x = 28.5

\mu = 30

\sigma = 3.5

n = 36

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (28.5 - 30) / 3.5 / \sqrt 36

= -2.57

Test statistic = -2.57

P(z < -2.57) = 0.0051

P-value = 0.0051

\alpha = 0.05

Z0.05 = -2.576

P-value > \alpha

Fail to reject the null hypothesis .

There is not enough evidence to suggest that the claim at \alpha = 0.05 .